OpenStudy (anonymous):

let f(x) = 2x +5 find the value of (f 0 f ^-1) (-1/2

OpenStudy (anonymous):

I dont understand the last part what does it say

OpenStudy (anonymous):

I think it's f composed of f inverse. \[(f \circ f^{-1})(-1/2) = f(f^{-1}(-1/2)) =\ ?\]

OpenStudy (anonymous):

yea.

OpenStudy (anonymous):

which is trivial to answer.

OpenStudy (anonymous):

Recall that a function and it's inverse undo each other.

OpenStudy (anonymous):

Did you figure it out?

OpenStudy (anonymous):

sort of not really. i knoe you plug in 1/2 for the inverse of 2x+5

OpenStudy (anonymous):

You really don't need to do anything of the sort. Recall that \(f(f^{-1}(x)) = x\) and therefore \((f\circ f^{-1})(-1/2) = f(f^{-1}(-1/2)) = -1/2\)

OpenStudy (anonymous):

The function and it's inverse undo each other. So whatever the inverse does to -1/2, the function will undo and you'll be right back at -1/2